A 4.62 gram sample of an unknown gas is found to occupy a volume of 1.50 L at a pressure of 989 mm Hg and a temperature of 30°C. The molecular weight of the unknown gas is  g/mol.
Answer #1

Finding the moles of gas by ideal gas law

PV= nRT

P= pressure = 989/760= 1.30 atm     

V= volume = 1.50 L

n= number of moles = ?

R= gas constant = 0.0821 LatmmolK

T = temperature = 30+273.15= 303.15 K

n= PV/RT= (1.30*1.50)/(0.0821*303.15)= 0.078 moles

molecular weight of gas = mass/moles = 4.62 g/ 0.078 mol = 58.96 g/mol